OpenStudy (anonymous):

Write the equation of the line in slope-intercept form given the point (4,3) and the slope of 1/2. Please dont explain to me just do it and show your work i will fan you and give you a medal Because I only have a few more minuts to turn i in please do alll the steps thanks

OpenStudy (anonymous):

@subway_surfers

OpenStudy (anonymous):

What good is it if you don't learn?

OpenStudy (anonymous):

i didnt ask for your opinian bloopman

OpenStudy (anonymous):

i just need to graduate highschool

OpenStudy (anonymous):

The slope intercept form is y=mx+b where m is the slope To find b, replace the x, y, and m 3=1/2(4)+b 3=2+b 1=b So, the equation is y=1/2x+1

OpenStudy (anonymous):

This is true, but I think you should take a peek at the terms of service before you make another post like this.

OpenStudy (anonymous):

sam stop skipping through the pages of your lesson and maybe you'll pass by yourself

OpenStudy (anonymous):

and put clothes on dude